How do you convert the following hyperbola, x^2/.2-(y+4)^2/1=1, to general form.
Answer by lwsshak3(11628)

How do you convert the following hyperbola, x^2/.2-(y+4)^2/1=1, to general form.
x^2/.2-(y+4)^2/1=1
LCD:2
10x^2-2(y+4)^2=2
10x^2-2(y^2+8y+16)=2
10x^2-2y^2-16y-32=2
10x^2-2y^2-16y=34
5x^2-y^2-8y=17